Please help me translate the following into a quadratic equation and solve it. Thanks!
The length of a rectangle garden is three times it width; if the area of the garden is 75 square meters; what are its dimensions?

Answer by scott8148(6628)

let x=width, so 3x=length

area equals length times width __ 75=3x(x) __ 75=3x^2

3x^2-75=0 __ dividing by 3 __ x^2-25=0 __ factoring __ (x+5)(x-5)=0

x+5=0 __ x=-5 __ negative value not realistic

x-5=0 __ x=5